OpenStudy (anonymous):

10. The moon is about 3.8 x’s 105 km from the Earth. Show the average orbital speed about Earth is 1026 m/s.

OpenStudy (anonymous):

just use use the equation gravitational force by earth = centrifugal force on moon. G*m(earth)*m(moon)/(r^2) = m(moon)*v^(2)/r where r denotes radius. and in this equation mass of moon will get cancelled on both sides so you just need to know the mass of earth.

OpenStudy (anonymous):

The equation for orbital speed is \[V = \sqrt{\frac{ Gm }{ r }}\]where r is the distance between the earth and the moon. m is the orbital mass of in question, in our case is the earth. G is the newtonian gravitationl constant if r = 3.8 x 10^7 m [distance from moon to earth] m = 5.9 x 10^24 kg [mass of earth] G = 6.67 x 10^-11 [constant] then you can prove that the speed of the earth is 1026 m/s

OpenStudy (anonymous):

@dibuaman can you show me where I'm going wrong with the equation you gave? I tried the method you have and keep getting V = 3237 m/s. here is the equation i am using with the numbers filled in: \[(6.67*10^{-11})*(5.97*10^{24})*(7.35*10^{22})/(38000000^{2}) = (7.35*10^{22})*v^{2}/38000000, \] where am I going wrong?